The analysis of a hydrocarbon revealed that it was 85.6281% C and 14.3719% H by mass. When1.8gofthegaswasstoredina1L flask at 93.4515◦C, it exerted a pressure of 489 Torr. What is the molecular formula of the hydrocarbon?
1
Expert's answer
2013-01-21T08:59:30-0500
Firstwe find the molar mass of the the gas
pV = m/M R T;
65 190*0.001 = 1.8 / M * 8.314 * 366.4515
M = 1.8 * 8.314 * 366.4515 / 65 190 * 0.001;
M = 84 g/mol
Then find the mass and amount of substances of each element
n = m/M
m(C) = 1.8 * 0.856281 = 1.54g; n (C) = 1.54 / 12 = 0.128 g/mol
m(H) = 1.8 * 0.143719 = 0.259g; n (H) = 0.259 / 1 = 0.259 g/mol
The amount of substances are correlated as:
n(H) : n(C) = 0.259 : 0.128 or 2 : 1
Knowing the the molar mass can construct a mathematical equation and to solve
it
12x + 1 * 2x= 84
14x = 84;
x = 84/14
x = 6
so the formula will be C6H12